# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v6

@@ Abstract
In this work we systematically review the recent advancements in software engineering with language models, covering 70+ models, 40+ evaluation tasks, 180+ datasets, and 900 related works. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also go beyond programming and review LLMs’ application in other software engineering activities including requirement engineering, testing, deployment, and operations in an endeavor to provide a global view of NLP in SE. We identify key challenges and potential future directions in this domain, and keep the survey open and updated on GitHub at https://github.com/codefuse-ai/Awesome-Code-LLM.
@@ 1 Introduction
The application of pretrained Transformers in software engineering, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and also goes beyond code processing (Arora et al., 2023; Feng et al., 2024; Ma et al., 2024). However, this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for SE in this work, covering 70+ models, 40+ downstream tasks, 180+ datasets, and 900 related works. We break down different categories of code language models, ranging from colossal models trained on general domains to tiny models trained specifically for code understanding or generation. We emphasize on the relations and differences between such models, and highlight the integration of code-specific features, such as abstract syntax trees or data flows, into language models, as well as the latest techniques adapted from NLP.
@@
Thus, building on these works, we endeavor to unite the perspectives from both communities, and accentuate the integration between NLP and SE throughout the work. We make the key observation that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2024; Luo et al., 2023), infilling objectives (Tay et al., 2023b; Li et al., 2023h; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023i), architectural improvements (Shazeer, 2019; Su et al., 2024; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements are providing real-world testbeds for these technologies and driving the development of LLMs forward into production. We believe a systematic review of these advancements would benefit both communities. Furthermore, all of the existing reviews focus on programming-related tasks, while we are the first to explicitly go beyond programming and provide a global view of LLM applications in the full life cycle of software development, covering distinct stages such as requirement engineering, deployment, and operations.
@@ 2 Downstream Tasks in Software Engineering
In this section, we first briefly introduce each of the traditional SE tasks and the application of pretrained language models in them in 2.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in 2.2 and investigate program synthesis in more detail in 2.3. Lastly, we also discuss the latest trend of repository-level coding in 2.4. In Appendix A, we list benchmarks for each downstream task.
@@ 2.1.1 Requirement Engineering
{forest} for tree= forked edges, draw, rounded corners, grow=east, anchor=base west, anchor=center, reversed=true, l sep=0.2cm, , where level=0font=, where level=1text width=4.9em,font=, where level=2text width=3.5em,font=, where level=3text width=10.5em,font=, where level=4text width=5.7em,font=, where level=5text width=10.8em,font=, for tree= if level=0l sep+=0.15cm, if level=1l sep+=0.15cm, if level=2l sep+=0.03cm, , [NLP in SE, fill=black, fill opacity=0.7, text=white [Requirement, fill=yellow, fill opacity=0.5 [Requirement Analysis, fill=yellow, fill opacity=0.5 [ ARSENAL (Ghosh et al., 2016), CAR (Ferrari et al., 2014), RETA (Arora et al., 2015a), NARCIA (Arora et al., 2015b), FENL (Bakar et al., 2016), SNACC (Zhang & El-Gohary, 2017), REGICE (Arora et al., 2017), AUR-BoW (Lu & Liang, 2017), ALERTme (Guzman et al., 2017), SAFE (Johann et al., 2017), OCLgen (Wang et al., 2018b), SEMIOS (Mezghani et al., 2018), Tiwari et al. (2019) , fill=yellow, fill opacity=0.5 [ , fill=yellow, fill opacity=0.5 [ Sainani et al. (2020), RE-BERT (de Araújo & Marcacini, 2021), White et al. (2023), Rodriguez et al. (2023), Jain et al. (2023), Arora et al. (2023), Ronanki et al. (2023), COREQQA (Abualhaija et al., 2023), CompliAT (Arora et al., 2024) , fill=yellow, fill opacity=0.5 ] ] ] ] [UI/UX
@@
Figure 3: NLP applications in the requirement engineering stage of SE. For each task the first column is non-neural methods; the second column is non-Transformer neural methods; the third column is Transformer-based methods.
- Requirement analysis refers to the process of extracting, summarizing, classifying, and validating software requirements. Most early works in this field rely on statistical NLP technologies such as POS (Part-of-Speech) tagging and dependency parsing, while more recent works utilize LLMs to classify, summarize, or extract requirement clauses. Zhao et al. (2020) provide a fine-grained survey on requirement analysis.
@@
Code Search, fill=Orange, fill opacity=0.2 [NL-to-Code
@@
Figure 4: NLP applications in the development stage of SE. For each task the first column is non-neural methods; the second column is non-Transformer neural methods; the third column is Transformer-based methods.
@@ 2.1.2 Development
Code Editing, fill=Orange, fill opacity=0.2 [Code
@@
Figure 5: NLP applications in the development stage of SE. For each task the first column is non-neural methods; the second column is non-Transformer neural methods; the third column is Transformer-based methods.
@@ 2.1.3 Testing and Analysis
{forest} for tree= forked edges, draw, rounded corners, grow=east, anchor=base west, anchor=center, reversed=true, l sep=0.2cm, , where level=0font=, where level=1text width=4.9em,font=, where level=2text width=3.5em,font=, where level=3text width=10.5em,font=, where level=4text width=5.7em,font=, where level=5text width=10.8em,font=, for tree= if level=0l sep+=0.15cm, if level=1l sep+=0.15cm, if level=2l sep+=0.03cm, , [NLP in SE, fill=black, fill opacity=0.7, text=white [Requirement, fill=yellow, fill opacity=0.5] [Development, fill=Orange, fill opacity=0.2] [Testing, fill=cyan, fill opacity=0.2 [Assertion Generation, fill=cyan, fill opacity=0.2 [ MeMo (Blasi et al., 2021), fill=cyan, fill opacity=0.2 [ Atlas (Watson et al., 2020), fill=cyan, fill opacity=0.2 [ Tufano et al. (2022a), TOGA (Dinella et al., 2022), FSLM (Bareiß et al., 2022), Yuan et al. (2023a), fill=cyan, fill opacity=0.2 ] ] ] ] [Mutant
@@
Figure 6: NLP applications in the testing & analysis stage of SE. For each task the first column is non-neural methods; the second column is non-Transformer neural methods; the third column is Transformer-based methods.
@@ 2.1.4 Deployment and Operations
Analysis, fill=cyan, fill opacity=0.2] [Deployment, fill=VioletRed, fill opacity=0.2 [Compiler
@@
Figure 7: NLP applications in the deployment & operations stage of SE. For each task the first column is non-neural methods; the second column is non-Transformer neural methods; the third column is Transformer-based methods.
@@ 8 Conclusion and Challenges
In this work, we systematically reviewed the history of code processing with pretrained Transformer language models, and highlighted their relations and comparisons to models pretrained on general domains. The advancement in code modeling generally follows the history course of NLP, evolving from SMT models, to NMT models, and then to finetuning pretrained Transformers and lastly to few-shot application of LLMs and even autonomous agents in real-world production. Unlike natural languages, the nature of code makes it easy to extract auxiliary information from alternative views, and to utilize interpreter and unit tests for automatic feedback.
@@
- Comprehensive benchmarks to push code LLMs to the next stage. The widely used HumanEval benchmark plays a key role in the evolution of Code LLMs. However, it is relatively small and its scoreboard has been manipulated to near perfect, which does not exactly reflect real-world behaviors. Many other benchmarks for Code LLMs have been proposed, but they are still not comprehensive enough to reflect production-level requirements. The community is eager for a new standard benchmark after HumanEval to further boost the progress of Code LLMs to the next stage.
@@
- Application of LLMs in more code downstream tasks. As we have pointed out in Section 2, current evaluation of LLMs’ coding capability in the NLP community is focused on program synthesis, and Figure 3-8 clearly show that tasks related to other activities in software engineering, such as software modeling and testing, have seen few application of LLMs. Besides, since the context window of LLMs are currently quite limited, generation tasks such as program synthesis and code translation are yet to be applied beyond method level. In Section 2.4, we have listed several works on repository-level code completion and temporal editing, and we believe the application of LLMs in more repository-level tasks will become a hot research topic in the future.
@@
- Building code LLM ecosystem for full-life-cycle of software development. While the academia have witnessed an abundance of code models, most have been deployed in the coding stage as IDE plugins while neglecting other stages in the life-cycle of software development. In Section 7.2 we mentioned several inspiring examples, and we are hoping to see more applications of code LMs throughout the full life-cycle of software development, from requirement engineering to DevOps, eventually leading to full-scale ecosystems like those around PyTorch (Paszke et al., 2019) and Hugging Face [^13] https://huggingface.co/.
@@
With the presentation of this survey, we hope to provide a global view of language models’ application in software engineering and connect the researches from the two communities. We believe the current surge of LLMs will be ultimately transformed into real world applications, and lead humanity into a brighter future.
@@ Acknowledgments
This work is supported by Ant Group.
@@ Appendix A Benchmarks for Downstrem Tasks
Task Date Benchmark Source Size Language Program Synthesis 2018-02 NL2Bash Lin et al. (2018b) 9305 Bash 2018-08 CONCODE Iyer et al. (2018) 104K Java 2019-10 JuICe Agashe et al. (2019) 1.5M/3725 Python 2021-05 APPS Hendrycks et al. (2021a) 10000 Python 2021-07 HumanEval Chen et al. (2021b) 164 Python 2021-08 MBPP Austin et al. (2021) 974 Python 2021-08 MathQA-Python Austin et al. (2021) 23914 Python 2021-08 PlotCoder Chen et al. (2021c) 40797 Python 2022-01 DSP Chandel et al. (2022) 1119 Python 2022-02 CodeContests Li et al. (2022h) 13610 C++, Python, Java 2022-03 MCoNaLa Wang et al. (2023i) 896 Python 2022-06 AixBench Hao et al. (2022) 336 Java 2022-10 MBXP Athiwaratkun et al. (2023) 12.4K 13 2022-10 Multilingual HumanEval Athiwaratkun et al. (2023) 1.9K 12 2022-10 MathQA-X Athiwaratkun et al. (2023) 5.6K Python, Java, JS 2022-11 ExeDS Huang et al. (2022b) 534 Python 2022-11 DS-1000 Lai et al. (2023) 1000 Python 2022-12 ODEX Wang et al. (2022h) 945 Python 2023-02 CoderEval Yu et al. (2023a) 460 Python, Java 2023-03 xCodeEval Khan et al. (2023) 5.5M 11 2023-03 HumanEval-X Zheng et al. (2023a) 820 Python, C++, Java, JS, Go 2023-06 StudentEval Babe et al. (2023) 1749 Python 2023-06 DotPrompts Agrawal et al. (2023) 10538 Java 2023-09 CodeApex Fu et al. (2023) 476 C++ 2023-09 VerilogEval Liu et al. (2023g) 8645/156 Verilog 2023-11 ML-Bench Liu et al. (2023l) 10040 Bash 2024-01 ParEval Nichols et al. (2024) 420 C++, HIP, CUDA 2024-04 PECC Haller et al. (2024) 2396 Python 2024-05 NaturalCodeBench Zhang et al. (2024b) 402 Python, Java 2024-05 Plot2Code Wu et al. (2024a) 132 Python 2024-05 MHPP Dai et al. (2024b) 140 Python
@@
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) 600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python 2024-03 ProCQA Li et al. (2024b) 5.2M C, C++, Java, Python, Ruby, Lisp, JS, C#, Go, Rust, PHP Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) 15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) 13948 2023-06 CMMLU Li et al. (2023b) 11528 2023-09 CodeApex Fu et al. (2023) 250 C++ 2024-01 CRUXEval Gu et al. (2024) 800 Python 2024-05 PythonIO Zhang et al. (2024d) 2650 Python Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) 300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) 513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023c) 744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) 21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python

### v7

@@ Abstract
In this work we systematically review the recent advancements in software engineering with language models, covering 70+ models, 40+ evaluation tasks, 180+ datasets, and 900 related works. Unlike previous works, we integrate software engineering (SE) with natural language processing (NLP) by discussing the perspectives of both sides: SE applies language models for development automation, while NLP adopts SE tasks for language model evaluation. We break down code processing models into general language models represented by the GPT family and specialized models that are specifically pretrained on code, often with tailored objectives. We discuss the relations and differences between these models, and highlight the historical transition of code modeling from statistical models and RNNs to pretrained Transformers and LLMs, which is exactly the same course that had been taken by NLP. We also go beyond programming and review LLMs’ application in other software engineering activities including requirement engineering, testing, deployment, and operations in an endeavor to provide a global view of NLP in SE, and identify key challenges and potential future directions in this domain. We keep the survey open and updated on GitHub at https://github.com/codefuse-ai/Awesome-Code-LLM.
@@ 1 Introduction
The application of pretrained Transformers in software engineering, however, can be traced back to dates before decoder-only autoregressive models became dominant (Feng et al., 2020; Liu et al., 2020), and also goes beyond code processing (Arora et al., 2023; Feng et al., 2024; Ma et al., 2024). However, this domain is yet to witness a comprehensive review. In an attempt to bridge the gap between natural language processing (NLP) community and software engineering (SE) community on the topic of language model applications, we undertake a panoramic survey of language models for SE in this work, covering 70+ models, 40+ downstream tasks, 180+ datasets, and 900 related works. Putting the emphasis on code language models, we analyze their development, discuss their differences from general language models, and highlight the integration of code-specific features such as abstract syntax trees or data flows, as well as the latest techniques adapted from NLP.
@@
Thus, building on these works, we endeavor to unite the perspectives from both communities, and highlight the relationship between NLP and SE: SE applies language models to various tasks for automation, while NLP adopts tasks from SE to evaluate language models. We observe that advanced topics from language modeling have been recently introduced into code processing, including instruction tuning (Honovich et al., 2023; Xu et al., 2024; Luo et al., 2023), infilling objectives (Tay et al., 2023b; Li et al., 2023h; Rozière et al., 2023), recontemplation of scaling laws (Hoffmann et al., 2022; Gunasekar et al., 2023; Li et al., 2023i), architectural improvements (Shazeer, 2019; Su et al., 2024; Dao et al., 2022), and autonomous agents (Qian et al., 2023; Hong et al., 2023), while in return SE requirements, represented by programming (Chen et al., 2021b), are providing real-world testbeds for these technologies and driving the development of LLMs forward into production and deployment. We believe a systematic review of these advancements would benefit both communities. Furthermore, unlike the existing reviews that focus on programming-related tasks, we are also the first to explicitly go beyond programming and provide a global view of LLM applications in the full life cycle of software development, covering distinct stages such as requirement engineering, deployment, and operations.
@@ 2 Downstream Tasks in Software Engineering
In this section, we first briefly introduce each of the traditional SE tasks and the application of pretrained language models in them in 2.1, and provide a comprehensive list of related works for each task. Then, we review the evaluation metrics in 2.2 and investigate program synthesis in more detail in 2.3. Lastly, we also discuss the latest trend of repository-level coding in 2.4. In Appendix A and B respectively, we list benchmarks for each downstream task and language models’ performance on them.
@@ 2.1.1 Requirement Engineering
{forest} for tree= forked edges, draw, rounded corners, grow=east, anchor=base west, anchor=center, reversed=true, l sep=0.2cm, , where level=0font=, where level=1text width=4.9em,font=, where level=2text width=3.5em,font=, where level=3text width=10.5em,font=, where level=4text width=5.7em,font=, where level=5text width=10.8em,font=, for tree= if level=0l sep+=0.15cm, if level=1l sep+=0.15cm, if level=2l sep+=0.03cm, , [NLP in SE, fill=black, fill opacity=0.7, text=white [Requirement, fill=yellow, fill opacity=0.5 [, draw=none, no edge [Non-neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Non-Transformer neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Transformer-based methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white] ] ] ] [Requirement Analysis, fill=yellow, fill opacity=0.5 [ ARSENAL (Ghosh et al., 2016), CAR (Ferrari et al., 2014), RETA (Arora et al., 2015a), NARCIA (Arora et al., 2015b), FENL (Bakar et al., 2016), SNACC (Zhang & El-Gohary, 2017), REGICE (Arora et al., 2017), AUR-BoW (Lu & Liang, 2017), ALERTme (Guzman et al., 2017), SAFE (Johann et al., 2017), OCLgen (Wang et al., 2018b), SEMIOS (Mezghani et al., 2018), Tiwari et al. (2019) , fill=yellow, fill opacity=0.5 [ , fill=yellow, fill opacity=0.5 [ Sainani et al. (2020), RE-BERT (de Araújo & Marcacini, 2021), White et al. (2023), Rodriguez et al. (2023), Jain et al. (2023), Arora et al. (2023), Ronanki et al. (2023), COREQQA (Abualhaija et al., 2023), CompliAT (Arora et al., 2024) , fill=yellow, fill opacity=0.5 ] ] ] ] [UI/UX
@@
Figure 3: NLP applications in the requirement engineering stage of SE.
- Requirement analysis refers to the process of extracting, summarizing, classifying, and validating software requirements. Most early works in this field rely on statistical NLP technologies such as POS (Part-of-Speech) tagging and dependency parsing to extract actions from requirements, while more recent works utilize LLMs to directly classify, summarize, or extract requirement clauses. Zhao et al. (2020) provide a fine-grained survey on requirement analysis.
@@
Code Search, fill=Orange, fill opacity=0.2 [, draw=none, no edge [Non-neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Non-Transformer neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Transformer-based methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white] ] ] ] [NL-to-Code
@@
Figure 4: NLP applications in the development stage of SE.
@@ 2.1.2 Development
Code Editing, fill=Orange, fill opacity=0.2 [, draw=none, no edge [Non-neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Non-Transformer neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Transformer-based methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white] ] ] ] [Code
@@
Figure 5: NLP applications in the development stage of SE.
@@ 2.1.3 Testing and Analysis
{forest} for tree= forked edges, draw, rounded corners, grow=east, anchor=base west, anchor=center, reversed=true, l sep=0.2cm, , where level=0font=, where level=1text width=4.9em,font=, where level=2text width=3.5em,font=, where level=3text width=10.5em,font=, where level=4text width=5.7em,font=, where level=5text width=10.8em,font=, for tree= if level=0l sep+=0.15cm, if level=1l sep+=0.15cm, if level=2l sep+=0.03cm, , [NLP in SE, fill=black, fill opacity=0.7, text=white [Requirement, fill=yellow, fill opacity=0.5] [Development, fill=Orange, fill opacity=0.2] [Testing, fill=cyan, fill opacity=0.2 [, draw=none, no edge [Non-neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Non-Transformer neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Transformer-based methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white] ] ] ] [Assertion Generation, fill=cyan, fill opacity=0.2 [ MeMo (Blasi et al., 2021), fill=cyan, fill opacity=0.2 [ Atlas (Watson et al., 2020), fill=cyan, fill opacity=0.2 [ Tufano et al. (2022a), TOGA (Dinella et al., 2022), FSLM (Bareiß et al., 2022), Yuan et al. (2023a), fill=cyan, fill opacity=0.2 ] ] ] ] [Mutant
@@
Figure 6: NLP applications in the testing & analysis stage of SE.
@@ 2.1.4 Deployment and Operations
Analysis, fill=cyan, fill opacity=0.2] [Deployment, fill=VioletRed, fill opacity=0.2 [, draw=none, no edge [Non-neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Non-Transformer neural methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white [Transformer-based methods, draw=none, no edge, fill=black, fill opacity=0.5, text=white] ] ] ] [Compiler
@@
Figure 7: NLP applications in the deployment & operations stage of SE.
@@ 8 Conclusion and Challenges
In this work, we systematically reviewed the history of pretrained Transformer language models in code processing and other software engineering tasks. The advancement in code modeling generally follows the history course of NLP, evolving from SMT models, to NMT models, and then to finetuning pretrained Transformers and lastly to few-shot application of LLMs and even autonomous agents in real-world production. Unlike natural languages, the nature of code makes it easy to extract auxiliary information from alternative views, and to utilize interpreter and unit tests for automatic feedback.
@@
- More comprehensive and challenging benchmarks. The widely used HumanEval benchmark plays a key role in the evolution of Code LLMs. However, it is relatively small and its scoreboard has been manipulated to near perfect, and the community is eager for a new standard to evaluate LLMs. HumanEval and other similar benchmarks focus on generating standalone Python functions from well-structured docstrings, which do not reflect real-world user behaviors. In real-world scenarios, software requirements are seldom condensed into a single docstring, and LLMs must learn to communicate effectively for clarifications when uncertain about the requirements (Wu & Fard, 2024). Also, standalone functions are hardly used in production, where software has complex dependencies within a repository. Thus, next-generation benchmarks should also take such cross-file context into account, and recent benchmarks such as SWE-bench (Jimenez et al., 2023) represent promising efforts in this direction. Contrary to the models’ perfect scores on HumanEval, these repository-level benchmarks expose LLMs’ limits in real-world applications, with the most recent SWE-agent resolving only 12.5% real-world GitHub issues in SWE-bench (Yang et al., 2024b).
- Evaluation and application of LLMs beyond traditional code generation. As we have pointed out in Section 2, current evaluation of LLMs’ coding capability in the NLP community is focused on code generation, while overlooking other activities in software engineering, such as software modeling and testing. From an application point-of-view, the most widespread application of LLMs in software engineering is IDE plugins, which provide developers with code suggestions, while other stages in software development that we mentioned in Section 2 are also largely overlooked. However, it should be also noted that LLMs trained on code are highly specialized to reasoning-heavy tasks such as programming and math, and non-coding tasks in SE - such as requirement elicitation, product management, and marketing - are probably better suited for general-domain models. Beyond the previously mentioned text-based evaluation and applications, the recent advancement of multimodel LLMs also created new opportunities, especially for UI design and other activities that involve visual input, leading to novel tasks such as visually grounded code generation (Li et al., 2024b; Wu et al., 2024a), webpage reverse engineering (Si et al., 2024; Laurençon et al., 2024), and layout design (Feng et al., 2023b).
@@
With the presentation of this survey, we hope to provide a global view of language models’ application in software engineering and connect the research from the two communities. We believe the current surge of LLMs will be ultimately transformed into real-world applications, and lead humanity into a brighter future.
@@ Acknowledgments
This work is supported by Ant Group. Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper. Only add this information once your submission is accepted and deanonymized.
@@ Appendix A Benchmarks for Downstrem Tasks
Task Date Benchmark Source Size Language Program Synthesis 2018-02 NL2Bash Lin et al. (2018b) 9305 Bash 2018-08 CONCODE Iyer et al. (2018) 104K Java 2019-10 JuICe Agashe et al. (2019) 1.5M/3725 Python 2021-05 APPS Hendrycks et al. (2021a) 10000 Python 2021-07 HumanEval Chen et al. (2021b) 164 Python 2021-08 MBPP Austin et al. (2021) 974 Python 2021-08 MathQA-Python Austin et al. (2021) 23914 Python 2021-08 PlotCoder Chen et al. (2021c) 40797 Python 2022-01 DSP Chandel et al. (2022) 1119 Python 2022-02 CodeContests Li et al. (2022h) 13610 C++, Python, Java 2022-03 MCoNaLa Wang et al. (2023i) 896 Python 2022-06 AixBench Hao et al. (2022) 336 Java 2022-10 MBXP Athiwaratkun et al. (2023) 12.4K 13 2022-10 Multilingual HumanEval Athiwaratkun et al. (2023) 1.9K 12 2022-10 MathQA-X Athiwaratkun et al. (2023) 5.6K Python, Java, JS 2022-11 ExeDS Huang et al. (2022b) 534 Python 2022-11 DS-1000 Lai et al. (2023) 1000 Python 2022-12 ODEX Wang et al. (2022h) 945 Python 2023-02 CoderEval Yu et al. (2023a) 460 Python, Java 2023-03 xCodeEval Khan et al. (2023) 5.5M 11 2023-03 HumanEval-X Zheng et al. (2023a) 820 Python, C++, Java, JS, Go 2023-06 StudentEval Babe et al. (2023) 1749 Python 2023-06 DotPrompts Agrawal et al. (2023) 10538 Java 2023-09 CodeApex Fu et al. (2023) 476 C++ 2023-09 VerilogEval Liu et al. (2023g) 8645/156 Verilog 2023-11 ML-Bench Liu et al. (2023l) 10040 Bash 2024-01 ParEval Nichols et al. (2024) 420 C++, HIP, CUDA 2024-04 MMCode Li et al. (2024b) 3548 Python 2024-04 PECC Haller et al. (2024) 2396 Python 2024-05 NaturalCodeBench Zhang et al. (2024b) 402 Python, Java 2024-05 Plot2Code Wu et al. (2024a) 132 Python 2024-05 MHPP Dai et al. (2024b) 140 Python
@@
Task Date Benchmark Source Size Language Code Retrieval 2018-03 StaQC Yao et al. (2018) 268K Python, SQL 2018-05 DeepCS Gu et al. (2018) 16M Java 2018-05 CoNaLa Yin et al. (2018) 600K/2.9K Python 2019-08 unnamed Li et al. (2019b) 287 Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M/99 Go, JS, Python, PHP, Java, Ruby 2020-02 CosBench Yan et al. (2020) 52 Java 2020-08 SO-DS Heyman & Cutsem (2020) 2.2K Python 2020-10 FB-Java Ling et al. (2021) 249K Java 2021-02 AdvTest Lu et al. (2021) 280K Python 2021-02 WebQueryTest Lu et al. (2021) 1K Python 2021-05 CoSQA Huang et al. (2021) 21K Python 2024-03 ProCQA Li et al. (2024c) 5.2M C, C++, Java, Python, Ruby, Lisp, JS, C#, Go, Rust, PHP Code Reasoning 2020-09 MMLU Hendrycks et al. (2021b) 15908 2021-09 CodeQA Liu & Wan (2021) 120K/70K Java/Python 2022-10 CS1QA Lee et al. (2022) 9237 2023-05 C-Eval Huang et al. (2023b) 13948 2023-06 CMMLU Li et al. (2023b) 11528 2023-09 CodeApex Fu et al. (2023) 250 C++ 2024-01 CRUXEval Gu et al. (2024) 800 Python 2024-05 PythonIO Zhang et al. (2024d) 2650 Python Type Inference 2019-12 TypeWriter OSS Pradel et al. (2020) 208K Python 2020-04 Typilus Allamanis et al. (2020) 252K Python 2020-04 LambdaNet Wei et al. (2020) 300 TypeScript 2021-04 ManyTypes4Py Mir et al. (2021) 869K Python 2022-10 ManyTypes4TypeScript Jesse & Devanbu (2022) 9.1M TypeScript 2023-02 TypeWeaver Yee & Guha (2023) 513 TypeScript 2023-03 BetterTypes4Py Wei et al. (2023) 608K Python 2023-03 InferTypes4Py Wei et al. (2023) 4.6K Python 2023-05 OpenTau Cassano et al. (2023c) 744 TypeScript Clone Detection / Code Search 2014-09 BigCloneBench Svajlenko et al. (2014) 6M Java 2014-09 POJ-104 Mou et al. (2016) 52K C, C++ 2019-05 unnamed Perez & Chiba (2019) 21K/24K Java/Python 2019-11 CLCDSA Nafi et al. (2019) 78K Java, C#, Python